When moving from infinity to towards a negative point chargem the electric potential

a. remains the same

b. undetermined

c. increases

d. decreases

Why do you say so? Explain.

1
Expert's answer
2021-04-28T07:23:57-0400

The potential at infinity is chosen to be zero.

Thus V for a point charge decreases with distance, whereas E for a point charge decreases with distance squared:

"E=\frac{F}{q}"

"E=\frac{kq}{r^2} =\frac{F}{q}"


Recall that the electric potential V is a scalar and has no direction, whereas the electric field E is a vector.

Then we say that option (d) is correct option

V is inversely proportional to distance r

But charge is negative (-Q)

"V= -\frac{kQ}{r}"

"V= -[\frac{kQ}{r}]"

Then option (d)is correct